### v2

# Sectorial covers over fanifolds

Abstract.
For the stopped Weinstein sector associated with any fanifold introduced in [GS23] by Gammage–Shende, we construct a Weinstein sectorial cover which allows us to describe homological mirror symmetry over the fanifold as an isomorphism of cosheaves of categories. In a special case, our Weinstein sectorial cover gives a lift of the open cover for the global skeleton of a very affine hypersurface computed in [GS22].
1. Introduction
The Fukaya category of a symplectic manifold is a highly nontrivial invariant. It is an $A_{\infty}$-category whose homotopy category has Lagrangians as objects and morphisms between two Lagrangians is the linear span of their intersection points up to suitable perturbation. The $A_{\infty}$-structure is controlled by counts of pseudoholomorphic disks with Lagrangian boundary conditions, which depend on the global geometry of the symplectic manifold. Usually, computing Fukaya categories is extremely hard because of this global nature. However, under suitable assumptions, Fukaya categories exhibit good local-to-global behaviors which allows us to reduce complicated analysis on the global geometry of symplectic manifolds to more tractable one.
For instance, via sectorial descent recently established in [GPS2] by Ganatra–Pardon–Shende, the wrapped Fukaya category $\mathcal{W}(X)$ of a Weinstein sector $X$ can be computed by gluing that of local pieces, provided its Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$. More precisely, the canonical functor from the homotopy colimit
$\displaystyle\operatorname{colim}_{\emptyset\neq I\subset\{1,\ldots,n\}}\mathcal{W}(\bigcap_{i\in I}X_{i})\to\mathcal{W}(X)$
induced by the pushforward functors along inclusions of Weinstein sectors from [GPS1] is a pretriangulated equivalence. In other words, wrapped Fukaya categories are cosheaves with respect to Weinstein sectorial covers. The result is valid also for partially wrapped Fukaya categories of stopped Weinstein sectors with mostly Legendrian stops away from boundaries.
Considering its computational power, given a Weinstein sector, it is natural to seek its Weinstein sectorial cover. Moreover, when working with homological mirror symmetry(HMS), one often needs such a cover to be compatible with that of its mirror. For instance, if the latter cover corresponds to an open cover of a relative skeleton, then a lift of it to a Weinstein sectorial cover will have the desired compatibility. So far, there are not so many interesting examples of Weinstein sectorial covers [CKL, GJ, GL22]. One reason for this might be easier access to other local-to-global behaviors [GS22, GS23, Lee16, MSZ, PS1, PS2, PS3], which requires us to compute at least locally a skeleton. Although they might work to solve a given problem, construction of Weinstein sectorial covers still should be a fundamental geometric question.
In this paper, we give a Weinstein sectorial cover of the stopped Weinstein sector associated with a fanifold introduced in [GS23] by Gammage–Shende to discuss HMS for more general spaces than toric stacks on the $B$-side. It is a stratified space obtained by gluing rational polyhedral fans of cones, which provides the organizing topological and discrete data for HMS at large volume. To a fanifold $\Phi$ they associated an algebraic space $\mathbf{T}(\Phi)$ obtained as the gluing of the toric varieties $T_{\Sigma}$ associated with the fans $\Sigma$. Based on an idea from SYZ fibrations, they constructed its mirror $\widetilde{\mathbf{W}}(\Phi)$ by inductive Weinstein handle attachments. The guiding principle behind their construction of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ was that the FLTZ Lagrangians $\mathbb{L}(\Sigma)$ and the projections $\mathbb{L}(\Sigma)\to\Sigma$ should glue to yield its relative skeleton $\widetilde{\mathbb{L}}(\Phi)$ and a map $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$.
As predicted by Gammage–Shende [GS23, Remark 4.5], the canonical lifts of the projections $\mathbb{L}(\Sigma)\to\Sigma$ glue to yield a version of $A$-side SYZ fibration $\widetilde{\mathbf{W}}(\Phi)\to\Phi$, after some modifications.
Theorem 1.1 ([Mor, Theorem 1.2]).
There is a filtered stratified fibration $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ restricting to $\pi$, which is homotopic to a filtered stratified integrable system with noncompact fibers. If the fan $\Sigma_{S}$ associated with any stratum $S\subset\Phi$ is proper, then the homotopy becomes trivial.
The lift $\bar{\pi}$ completes a global combinatorial duality for the mirror pair $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ over $\Phi$. Moreover, it enables us to lift any cover of $\Phi$ to $\widetilde{\mathbf{W}}(\Phi)$. Inspired by the idea in [GPS2, Example 1.34] and using this aspect of $\bar{\pi}$, after some modification, we construct a Weinstein sectorial cover of $\widetilde{\mathbf{W}}(\Phi)$ with the desired compatibility in Section $4$.
Theorem 1.2.
There exists a Weinstein sectorial cover of $\widetilde{\mathbf{W}}(\Phi)$. Moreover, it restricts to an open cover of the relative skeleton $\widetilde{\mathbb{L}}(\Phi)$ compatible with a cover of $\mathbf{T}(\Phi)$.
When $\Phi$ is the fanifold from [GS23, Example 4.23], $\widetilde{\mathbb{L}}(\Phi)$ coincides with the skeleton $\operatorname{Core}(H)$ [GS22, Theorem 6.2.4] of a very affine hypersurface $H$. In Section $5$, we show that our cover gives a lift of the open cover [GS22, Corollary 4.3.2] of $\operatorname{Core}(H)$, giving an affirmative answer to the conjecture raised in [GS22, Section 2.4] and making the discussion of [GS22, Section 2.1] completely rigorous.
Via sectorial descent we obtain a description of HMS for $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ as an isomorphism of cosheaves of categories. See Section $4.3$ for the precise definitions of the associated cosheaves. The isomorphism can be regarded as a dual to [GS23, Theorem 5.3] and encodes the gluing of coherent constructible correspondence [Kuw20] for local pieces.
Corollary 1.3 (Theorem 4.17).
There is an isomorphism
$\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$
of cosheaves of categories over $\Phi$ whose global section yields an equivalence
$\operatorname{Fuk}(\widetilde{\mathbf{W}}(\Phi))\simeq\operatorname{Coh}(\mathbf{T}(\Phi)).$
Our construction of the cover is close to the more general strategy suggested in [BC23] by Bai–Côté. When $X=T^{*}M$, they triangulate $M$ and cover $X$ by the cotangent bundles $T^{*}\operatorname{star}(v_{\alpha})$ of the stars of the vertices $v_{\alpha}$. When $X$ is a polarizable Weinstein manifold, they deform $X$ to another Weinstein manifold with arborial skeleton which by [BC23, Section 4.2] admits a canonical Whitney stratification, in particular, a triangulation due to Goresky. Cover the deformation of $X$ by arboreal thickenings of the stars of the vertices. Unfortunately, there are obstructions to making their strategy rigorous which involves developing a good theory of arboreal sectors.
Another general strategy might come from simplicial decompositions introduced in [Asp23] by Asplund. By [Asp23, Theorem 1.4] a simplicial decomposition of a Weinstein manifold $X$ bijectively corresponds to its sectorial cover which is good in the sense of [Asp23, Section 1]. Since the convex completion of $\widetilde{\mathbf{W}}(\Phi)$ is an ordinary Weinstein manifold, by [Asp23, Lemma 3.13] one could refine our Weinstein sectorial cover to be a good sectorial cover of the convex completion. It would be an interesting problem to find a simplicial decomposition of a Weinstein manifold corresponding to a Weinstein sectorial cover and compatible with the cover of its given mirror.
Acknowledgements.
The author was supported by SISSA PhD scholarships in Mathematics. This work was partially supported by JSPS KAKENHI Grant Number JP23KJ0341.
2. Gluing of stopped Weinstein sectors
In this section, we review some basic properties of Liouville sectors introduced in [GPS1, Section 2] by Ganatra–Pardon–Shende and how to glue them together. Throughout the paper, we work with Liouville manifolds of finite type so that any Liouville manifold $X$ will be the completion of some Liouville domain $X_{0}$. Then its skeleton $\operatorname{Core}(X)$ will always be compact. We denote by $\partial_{\infty}X_{0}$ the actual boundary of $X_{0}$, which can be identified with the ideal boundary $\partial_{\infty}X$ of $X$ to preserve the symbols $\partial X_{0},\partial X$.
2.1. Stopped Weinstein sectors
All Liouville manifolds of our interest are Weinstein. Namely, the Liouville vector field $Z$ of each $X$ is gradient-like for a Morse–Bott function $\phi\colon X\to\mathbb{R}$ which is constant on the cylindrical end $(\mathbb{R}_{t\geq 0}\times\partial_{\infty}X,e^{t}(\lambda|_{\partial_{\infty}X}))$. Hence $\operatorname{Core}(X)$ will also be isotropic by [CE12, Lemma 11.13(a)].
Definition 2.1 ([GPS1, Definition 2.4]).
A Liouville sector is a Liouville manifold-with-boundary satisfying the following equivalent conditions:
  • For some $\alpha>0$ there exists an $\alpha$-defining function $I\colon\partial X\to\mathbb{R}$ with $ZI=\alpha I$ near infinity and $dI|_{C}>0$ for the characteristic foliation of the hypersurface $\partial X\subset X$, where $C$ is oriented so that $\omega(C,N)>0$ for any inward pointing vector $N$.
  • The boundary of the contact hypersurface $\partial_{\infty}X\subset X$ is convex and there is a diffeomorphism $\partial X\cong F\times\mathbb{R}$ sending $C$ to the foliation of $F\times\mathbb{R}$ by leaves $\{pt\}\times\mathbb{R}$. We call the symplectic reduction $F$, which is a Liouville manifold, the symplectic boundary of $X$.
An inclusion $i\colon X\hookrightarrow X^{\prime}$ of Liouville sectors is a proper map which is a diffeomorphism onto its image, satisfying $i^{*}\lambda^{\prime}=\lambda+df$ for compactly supported $f$. A trivial inclusion of Liouville sectors is one for which $i(X)$ may be deformed into $X^{\prime}$ through Liouville sectors included in $X$.
Remark 2.2.
The inequality $dI|_{C}>0$ is equivalent to the Hamiltonian vector field $X_{I}$ being outward pointing along $\partial X$. Small deformations of $X$ within the class of Liouville manifolds-with-boundary are again Liouville sectors.
Definition 2.3 ([GPS1, Remark 2.8]).
An open Liouville sector is a pair $(X,\partial_{\infty}X)$ of an exact symplectic manifold $X$ and a contact manifold $\partial_{\infty}X$ together with a germ near +$\infty$+ ∞ of codimension 00 embedding of the symplectization $S\partial_{\infty}X$ into $X$. Here, the embedding strictly respects Liouville forms and the pair $(X,\partial_{\infty}X)$ is exhausted by Liouville sectors. Being exhausted by Liouville sectors means that any subset of $X$, which away from a compact subset of $X$ is the cone over a compact subset of $\partial_{\infty}X$, is contained in some Liouville sector $X_{0}\subset X$ with $\partial_{\infty}X_{0}\subset\partial_{\infty}X$.
Example 2.4.
For any manifold $M$ the cotangent bundle $T^{*}M$ is an open Liouville sector. An exhaustion is given by the family of Liouville subsectors $T^{*}M_{0}$ for compact codimension 00 submanifolds-with-boundary $M_{0}\subset M$. For any Liouville sector $X$ its interior $X^{\circ}=X\setminus\partial X$ is an open Liouville sector.
For a Liouville sector $X$ we denote by $\mathcal{W}(X)$ the wrapped Fukaya category [GPS1, Definition 3.36]. It is an $A_{\infty}$-category whose objects are exact cylindrical at infinity Lagrangians inside $X$ and whose morphisms are calculated by wrapping Lagrangians. According to [GPS1, Section 3.6], any inclusion $X\hookrightarrow X^{\prime}$ of Liouville sectors induces the pushforward functor $\mathcal{W}(X)\to\mathcal{W}(X^{\prime})$ which makes the assignment $X\mapsto\mathcal{W}(X)$ covariantly functorial. We use the same symbols to denote the counterparts of open Liouville sectors.
Wrapped Fukaya categories are invariant under deformations in the following sense.
Lemma 2.5 ([GPS2, Lemma 3.4]).
Along a trivial inclusion $X\hookrightarrow X^{\prime}$ of Liouville sectors, the pushforward functor defines a canonical quasi-equivalence $\mathcal{W}(X)=\mathcal{W}(X^{\prime})$.
Definition 2.6 ([GPS2, Section 1.1]).
A stop $\mathfrak{f}$ of a Liouville sector $X$ is a closed subset of $(\partial_{\infty}X)^{\circ}=\partial_{\infty}X\setminus\partial\partial_{\infty}X$. We call the pair $(X,\mathfrak{f})$ a stopped Liouville sector. An inclusion $(X,\mathfrak{f})\hookrightarrow(X^{\prime},\mathfrak{f}^{\prime})$ of stopped Liouville sectors is an inclusion of Liouville sectors $X\hookrightarrow X^{\prime}$ with $\mathfrak{f}^{\prime}\cap(\partial_{\infty}X)^{\circ}\subset\mathfrak{f}$.
When $\mathfrak{f}=\emptyset$, the above definition specializes to Definition 2.1. For a stopped Liouville sector $(X,\mathfrak{f})$ we denote by $\mathcal{W}(X,\mathfrak{f})$ the partially wrapped Fukaya category [GPS2, definition 2.31]. It is an $A_{\infty}$-category whose objects are exact cylindrical at infinity Lagrangians inside $X$ disjoint at infinity from $\mathfrak{f}$ and whose morphisms are calculated by wrapping Lagrangians in $(\partial_{\infty}X)^{\circ}\setminus\mathfrak{f}$. Similar to wrapped Fukaya categories, any inclusion $(X,\mathfrak{f})\hookrightarrow(X^{\prime},\mathfrak{f}^{\prime})$ of stopped Liouville sectors induces the pushforward functor $\mathcal{W}(X,\mathfrak{f})\to\mathcal{W}(X^{\prime},\mathfrak{f}^{\prime})$ which makes the assignment $(X,\mathfrak{f})\mapsto\mathcal{W}(X,\mathfrak{f})$ covariantly functorial.
Although the partially wrapped Fukaya category is defined for an arbitrary stopped Liouville sector $(X,\mathfrak{f})$, the results in [GPS2] are generally sharpest when $X$ is Weinstein and $\mathfrak{f}$ is mostly Legendrian.
Definition 2.7 ([GPS2, Definition 1.7]).
A closed subset $\mathfrak{f}$ of a contact manifold $Y^{2n-1}$ is mostly Legendrian if it admits a decomposition $\mathfrak{f}=\mathfrak{f}^{\text{subcrit}}\cup\mathfrak{f}^{\text{crit}}$ for which $\mathfrak{f}^{\text{subcrit}}$ is closed and is contained in the smooth image of a second countable manifold of dimension $<n-1$, and $\mathfrak{f}^{\text{crit}}\subset Y\setminus\mathfrak{f}^{\text{subcrit}}$ is a Legendrian submanifold. The notion of a mostly Lagrangian closed subset of a symplectic manifold is defined analogously.
2.2. Weinstein pairs
Lemma 2.8 ([GPS1, Lemma 2.13]).
Let $\bar{X}_{0}$ be a Liouville domain and $A\subset\partial_{\infty}\bar{X}_{0}$ a codimension 00 submanifold-with-boundary for which there exists a function $I\colon A\to\mathbb{R}$ with $R_{\lambda}I>0$ such that the contact vector field $V_{I}$ is outward pointing along $\partial A$. Then $X=\bar{X}\setminus(A^{\circ}\times\mathbb{R}_{>0})$ is a Liouvile sector where $\bar{X}$ denotes the conic completion of $\bar{X}_{0}$.
Definition 2.9 ([GPS1, Definition 2.14]).
A sutured Liouville domain $(\bar{X}_{0},F_{0})$ is a Liouville domain $\bar{X}_{0}$ together with a codimension $1$ submanifold-with-boundary $F_{0}\subset\partial_{\infty}\bar{X}_{0}$ and a contact form $\lambda$ defined over a neighborhood $\operatorname{Nbd}_{\partial_{\infty}\bar{X}_{0}}(F_{0})$ of $F_{0}$ in $\partial_{\infty}\bar{X}_{0}$ such that $(F_{0},\lambda|_{F_{0}})$ is a Liouville domain. A sutured Liouville manifold is a Liouville manifold $\bar{X}$ together with a codimension $1$ submanifold-with-boundary $F_{0}\subset\partial_{\infty}\bar{X}$ and a contact form $\lambda$ defined over a neighborhood $\operatorname{Nbd}_{\partial_{\infty}\bar{X}}(F_{0})$ of $F_{0}$ in $\partial_{\infty}\bar{X}$ such that $(F_{0},\lambda|_{F_{0}})$ is a Liouville domain.
Remark 2.10.
Any sutured Liouville domain $(\bar{X}_{0},F_{0})$ defines a Weinstein pair in the sense of [Eli, Section 2] if $\lambda$ extends to $\bar{X}_{0}$ and there exists a function $\phi\colon\bar{X}_{0}\to\mathbb{R}$ which makes the triples $(\bar{X}_{0},\lambda,\phi)$ and $(F_{0},\lambda|_{F_{0}},\phi|_{F_{0}})$ Weinstein.
Definition 2.11 ([GPS1, Remark 2.16]).
Given a sutured Liouville domain $(\bar{X}_{0},F_{0})$ let $\mathbb{L}_{0}\subset\bar{X}_{0}$ be the locus of points which do not escape to the complement of $\operatorname{Core}(F_{0})$ at infinity. We call $\mathbb{L}_{0}$ the relative skeleton of $(\bar{X}_{0},F_{0})$ and denote it by $\operatorname{Core}(\bar{X}_{0},F_{0})$. Analogous terminology applies to sutured Liouville manifolds.
Remark 2.12.
The locus $\mathbb{L}\subset\bar{X}$ is necessarily noncompact unless $F_{0}$ is empty. When both $\bar{X},F_{0}$ are Weinstein, $\mathbb{L}$ is a singular isotropic spine. In particular, $\bar{X}$ deforms down to a small regular neighborhood of $\mathbb{L}$.
Given a sutured Liouville domain $(\bar{X}_{0},F_{0})$, the Reeb vector field $R_{\lambda}$ is transverse to $F_{0}$ as $d\lambda|_{F_{0}}$ is symplectic and determines a local coordinate charts $F_{0}\times\mathbb{R}_{|t|\leq\epsilon}\hookrightarrow\partial_{\infty}\bar{X}_{0}$ in which we have $\lambda=\lambda|_{F_{0}}+dt$. The contact vector field $V_{t}$ associated with the function $t$ is given by $Z|_{F_{0}}+t\partial_{t}$ which is outward pointing along $\partial(F_{0}\times\mathbb{R}_{|t|\leq\epsilon})$. Hence $(\bar{X}_{0},F_{0})$ determines a codimension 00 submanifold-with-boundary $A=F_{0}\times\mathbb{R}_{|t|\leq\epsilon}$ satisfying the hypotheses of Lemma 2.8. In particular, $(\bar{X}_{0},F_{0})$ gives rise to a Liouville sector.
Conversely, any Liouville sector arises, up to deformation, from a unique in the homotopical sense sutured Liouville domain. The following lemma plays a role to prove Lemma 2.17 below.
Lemma 2.13 ([GPS1, Lemma 2.23]).
Any pair $(\bar{X}_{0},A)$ satisfying the hypotheses of Lemma 2.8 arises, up to deformation, from a unique in the homotopical sense sutured Liouville domain.
Lemma 2.14 ([GPS1, Proposition 2.25]).
Let $X$ be a Liouville sector. Equip $\mathbb{C}$ the standard symplectic form $\omega_{\mathbb{C}}=dx\wedge dy$. Let $\lambda^{\alpha}_{\mathbb{C}}$ be the Liouville form associated with a Liouville vector field $Z^{\alpha}_{\mathbb{C}}=(1-\alpha)x\partial_{x}+\alpha y\partial_{y}$. Then any $\alpha$-defining function $I\colon\operatorname{Nbd}^{Z}_{X}(\partial X)\to\mathbb{R}$ extends to a unique identification
$(X,\lambda)=(F\times\mathbb{C}_{\operatorname{Re}\geq 0},\lambda_{F}+\lambda^{\alpha}_{\mathbb{C}}+df)$
valid over some neighborhoods of the respective boundaries, where $I=y$ is the imaginary part of the coordinate on $\mathbb{C}$, $(F,\lambda_{F})$ is a Liouville manifold and $f\colon F\times\mathbb{C}_{\operatorname{Re}\geq 0}\to\mathbb{R}$ satisfies the following conditions:
  • $f$ is supported inside $F_{0}\times\mathbb{C}$ for some Liouville domain $F_{0}\subset F$.
  • $f$ coincides with some $f_{\pm\infty}\colon F\to\mathbb{R}$ for sufficiently large $|I|$.
Definition 2.15 ([GPS1, Section 2.7]).
Let $X$ be a Liouville sector. The convex completion is the pair $(\bar{X},\lambda_{\bar{X}})$ of an exact symplectic manifold $\bar{X}=X\cup_{\operatorname{Nbd}\partial(F\times\mathbb{C}_{\operatorname{Re}\geq 0})}\operatorname{Nbd}(F\times\mathbb{C}_{\operatorname{Re}\leq 0})$ and a $1$-form $\lambda_{\bar{X}}=\lambda_{F}+\lambda^{1/2}_{\mathbb{C}}+d\tilde{f}$ for an extension $\tilde{f}\colon\operatorname{Nbd}(F\times\mathbb{C}_{\operatorname{Re}\leq 0})\to\mathbb{R}$ of $f\colon\operatorname{Nbd}\partial(F\times\mathbb{C}_{\operatorname{Re}\geq 0})\to\mathbb{R}$.
Lemma 2.16 ([GPS1, Lemma 2.31]).
If the extension $\tilde{f}\colon F\times\mathbb{C}_{\operatorname{Re}<\epsilon}\to\mathbb{R}$ is bounded uniformly in $C^{\infty}$ and satisfies the two bulleted properties from Lemma 2.14, then the convex completion $(\bar{X},\lambda_{\bar{X}})$ defines a Liouville manifold.
Lemma 2.17 ([GPS1, Lemma 2.32]).
Any Liouville sector arises, up to deformation, from a unique in the homotopical sense sutured Liouville manifold. Moreover, the convex completion of the Liouville sector $X$ associated with a sutured Liouville manifold $(\bar{X},F_{0})$ coincides with $\bar{X}$, and the inclusion $X\hookrightarrow\bar{X}$ is the obvious one.
2.3. Gluing of Weinstein pairs
Given a Weinstein pair $(\bar{X}_{0},F_{0})$, by [Eli, Proposition 2.9] one can always modify the Liouville form $\lambda$ for $\omega$ and the Morse–Bott function $\phi_{F_{0}}$ to be adjusted in the following sense.
Definition 2.18 ([Eli, Section 2]).
Let $(\bar{X}_{0},F_{0})$ be a Weinstein pair, $\phi_{F_{0}}\colon F_{0}\to\mathbb{R}$ a Morse–Bott function for which $Z_{F_{0}}=(\omega|_{F_{0}})^{\#}\lambda|_{F_{0}}$ is gradient-like and $U_{\epsilon}(F_{0})\subset\partial_{\infty}\bar{X}_{0}$ the contact surrounding of $F_{0}$. We call a pair $(\lambda_{0},\phi_{0})$ of a Liouville form $\lambda_{0}$ for $\omega$ and a smooth function $\phi_{0}\colon\bar{X}_{0}\to\mathbb{R}$ adjusted to $(\bar{X}_{0},F_{0})$ if
  • $Z_{0}=\omega^{\#}\lambda_{0}$ is tangent to $\partial_{\infty}\bar{X}_{0}$ on $U_{\epsilon}(F_{0})$ and transverse to $\partial_{\infty}\bar{X}_{0}$ elsewhere;
  • $Z_{0}|_{U_{\epsilon}(F_{0})}=Z_{F_{0}}+u\partial_{u}$;
  • the attractor $\bigcap_{t>0}Z^{-t}_{0}(\bar{X}_{0})$ coincides with $\operatorname{Core}(\bar{X}_{0},F_{0})$;
  • the function $\phi_{0}$ is Morse–Bott for which $Z_{0}=(\omega)^{\#}\lambda_{0}$ is gradient-like satisfying $\phi_{0}|_{U_{\epsilon}(F_{0})}=\phi_{F_{0}}+\frac{1}{2}u^{2}$ and whose critical values are not more than $\phi_{0}|_{\partial U_{\epsilon}(F_{0})}$.
Definition 2.19 ([Eli, Section 3.1]).
Let $(\bar{X}_{0},\lambda,\phi)$ be a Weinstein domain. A splitting for $(\bar{X}_{0},\lambda,\phi)$ is a hypersurface $(P,\partial_{\infty}P)\subset(\bar{X}_{0},\partial_{\infty}\bar{X}_{0})$ satisfying the following conditions:
  • $\partial_{\infty}P$ and $P$ respectively split $\partial_{\infty}\bar{X}_{0}$ and $\bar{X}_{0}$ into two parts $\partial_{\infty}\bar{X}_{0}=Y_{-}\cup Y_{+}$ with $\partial_{\infty}Y_{-}=\partial_{\infty}Y_{+}=Y_{-}\cap Y_{+}=\partial_{\infty}P$ and $\bar{X}_{0}=\bar{X}_{0-}\cup\bar{X}_{0+}$ with $\partial\bar{X}_{0-}=P\cup Y_{-},\partial\bar{X}_{0+}=P\cup Y_{+},\bar{X}_{0-}\cap\bar{X}_{0+}=P$.
  • The Liouville vector field $Z=\omega^{\#}\lambda$ is tangent to $P$.
  • There exists a hypersurface $(F_{0},\partial_{\infty}F_{0})\subset(P,\partial_{\infty}P)$ which is Weinstein for $\lambda|_{F_{0}}$, tangent to $Z$ and intersects transversely all leaves of the characteristic foliation of $P$.
The above hypersurface $(F_{0},\partial_{\infty}F_{0})$ is called the Weinstein soul for the splitting hypersurface $P$. Due to Lemma 2.21 below, $P$ is contactomorphic to the contact surrounding $U_{\epsilon}(F_{0})$ of its Weinstein soul.
Definition 2.20 ([Eli, Section 2]).
Let $F_{0}$ be a closed hypersurface in a $(2n-1)$-dimensional manifold and $\xi$ a germ of a contact structure along $F_{0}$ which admits a transverse contact vector field $V$. The invariant extension of the germ $\xi$ is the canonical extension $\hat{\xi}$ on $F_{0}\times\mathbb{R}$, which is invariant with respect to translations along the second factor and whose germ along any slice $F_{0}\times\{t\},t\in\mathbb{R}$ is isomorphic to $\xi$.
Lemma 2.21 ([Eli, Lemma 2.6]).
Let $F_{0}$ be a closed $(2n-2)$-dimensional manifold and $\xi$ a contact structure on $P=F_{0}\times[0,\infty)$ which admits a contact vector field $V$ inward transverse to $F_{0}\times\{0\}$ such that its trajectories intersecting $F_{0}\times\{0\}$ fill the whole manifold $P$. Then $(P,\xi)$ is contactomorphic to $(F_{0}\times[0,\infty),\hat{\xi})$, where $\hat{\xi}$ is the invariant extension of the germ of $\xi$ along $F_{0}\times\{0\}$. Moreover, for any compact set $C\subset P$ with $F_{0}\times\{0\}\subset\operatorname{Int}C$ there exists a contactomorphism $(P,\xi)\to(F_{0}\times[0,\infty),\hat{\xi})$, which equals the identity on $F_{0}\times\{0\}$ and sends $V|_{C}$ to the vector field $\partial_{t}$.
Definition 2.22 ([Eli, Section 3.1]).
Let $(\bar{X}_{0},F_{0}),(\bar{X}^{\prime}_{0},F^{\prime}_{0})$ be Weinstein pairs with adjusted Liouville forms and Morse–Bott functions $(\lambda,\phi),(\lambda^{\prime},\phi^{\prime})$. Suppose that there is an isomorphism $(F_{0},\lambda|_{F_{0}},\phi|_{F_{0}})\cong(F^{\prime}_{0},\lambda^{\prime}|_{F^{\prime}_{0}},\phi^{\prime}|_{F^{\prime}_{0}})$ of Weinstein manifolds. Extend it to their contact surroundings $U_{\epsilon}(F_{0})\cong U_{\epsilon}(F^{\prime}_{0})$. Then the gluing of $(\bar{X}_{0},F_{0}),(\bar{X}^{\prime}_{0},F^{\prime}_{0})$ is given by
$\bar{X}_{0}\#_{F_{0}\cong F^{\prime}_{0}}\bar{X}^{\prime}_{0}=\bar{X}_{0}\cup_{U_{\epsilon}(F_{0})\cong U_{\epsilon}(F^{\prime}_{0})}\bar{X}^{\prime}_{0}$
with glued Liouville form and Morse–Bott function.
By definition we have
$\operatorname{Core}(\bar{X}_{0}\#_{F_{0}\cong F^{\prime}_{0}}\bar{X}^{\prime}_{0})=\operatorname{Core}(\bar{X}_{0},F_{0})\cup_{\operatorname{Core}(F_{0})\cong\operatorname{Core}(F^{\prime}_{0})}\operatorname{Core}(\bar{X}^{\prime}_{0},F^{\prime}_{0}).$
3. Large volume limit fibrations over fanifolds
In this section, after the construction of the stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ associated with a fanifold $\Phi$ introduced in [GS23] by Gammage–Shende, we review that of the filtered stratified fibration $\bar{\pi}\colon\widetilde{\mathbf{W}}(\Phi)\to\Phi$ from Theorem 1.1 restricting to that $\pi\colon\widetilde{\mathbb{L}}(\Phi)\to\Phi$ of the relative skeleton $\widetilde{\mathbb{L}}(\Phi)$. The stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ is obtained by inductively attaching products of the cotangent bundles of real tori and strata of $\Phi$. Each step requires us to modify Weinstein structures near gluing regions.
3.1. Fanifolds
Throughout the paper, we will work with only those stratified spaces $\Phi$ which satisfy the following conditions:
  • (i)
    $\Phi$ has finitely many strata.
  • (ii)
    $\Phi$ is conical in the complement of a compact subset.
  • (iii)
    $\Phi$ is given as a germ of a closed subset in an ambient manifold $\mathcal{M}$.
  • (iv)
    The strata of $\Phi$ are smooth submanifolds of $\mathcal{M}$.
  • (v)
    The strata of $\Phi$ are contractible.
We will express properties of $\Phi$ in terms of $\mathcal{M}$ as long as they only depend on the germ of $\Phi$. Taking the normal cone $C_{S}\Phi\subset T_{S}\mathcal{M}$ for each stratum $S\subset\Phi$, one obtains a stratification on $C_{S}\Phi$ induced by that of a sufficiently small tubular neighborhood $T_{S}\mathcal{M}\to\mathcal{M}$.
Definition 3.1.
The stratified space $\Phi$ is smoothly normally conical if for each stratum $S\subset\Phi$ some choice of tubular neighborhood $T_{S}\mathcal{M}\to\mathcal{M}$ induces locally near $S$ a stratified diffeomorphism $C_{S}\Phi\to\Phi$, which in turn induces the identity $C_{S}\Phi\to C_{S}\Phi$.
Definition 3.2.
We write $\operatorname{Fan}^{\twoheadrightarrow}$ for the category whose objects are pairs $(M,\Sigma)$ of a laticce $M$ and a stratified space $\Sigma$ by finitely many rational polyhedral cones in $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$. For any $(M,\Sigma),(M^{\prime},\Sigma^{\prime})\in\operatorname{Fan}^{\twoheadrightarrow}$ a morphism $(M,\Sigma)\to(M^{\prime},\Sigma^{\prime})$ is given by the data of a cone $\sigma\in\Sigma$ and an isomorphism $M/\langle\sigma\rangle\cong M^{\prime}$ such that $\Sigma^{\prime}=\Sigma/\sigma=\{\tau/\langle\sigma\rangle\subset M_{\mathbb{R}}/\langle\sigma\rangle\ |\ \tau\in\Sigma,\sigma\subset\bar{\tau}\}$. We denote by $\operatorname{Fan}^{\twoheadrightarrow}_{\Sigma/}$ for $(M,\Sigma)\in\operatorname{Fan}^{\twoheadrightarrow}$ the full subcategory of objects $(M^{\prime},\Sigma^{\prime})$ with $\Sigma^{\prime}=\Sigma/\sigma$ for some $\sigma\in\Sigma$.
Definition 3.3 ([GS23, Definition 2.4]).
A fanifold is a smoothly normally conical stratified space $\Phi\subset\mathcal{M}$ equipped with the following data:
  • A functor $\operatorname{Exit}(\Phi)\to\operatorname{Fan}^{\twoheadrightarrow}$ from the exit path category of $\Phi$ whose value on each stratum $S$ is a pair $(M_{S},\Sigma_{S})$ of a lattice $M_{S}$ and a rational polyhedral fan $\Sigma_{S}\subset M_{S,\mathbb{R}}$ called the associated normal fan.
  • For each stratum $S\subset\Phi$ a trivialization $\phi_{S}\colon T_{S}\mathcal{M}\cong M_{S,\mathbb{R}}$ of the normal bundle carrying the induced stratification on $C_{S}\Phi$ to the standard stratification induced by $\Sigma_{S}$.
These data must make the diagram
[…]
commute for any stratum $S^{\prime}$ of the induced stratification on $\operatorname{Nbd}(S)$, where the left vertical arrow is the quotient by the span of $S^{\prime}$. The right vertical arrow corresponds to the map $M_{S}\to M_{S^{\prime}}$ on lattices.
Remark 3.4.
One can identify the exit path category $\operatorname{Exit}(\Sigma)$ with $\operatorname{Fan}^{\twoheadrightarrow}_{\Sigma/}$ as a poset via
$\sigma\mapsto[\Sigma\mapsto\Sigma/\sigma].$
In addition, the normal geometry to $\sigma$ is the geometry of $\Sigma/\sigma$. Due to the condition (v), $\operatorname{Exit}_{S}(\Phi)$ is equivalent to the poset $\operatorname{Exit}(\Sigma_{S})$, the full subcategory of exit paths starting at $S$ contained inside a sufficiently small neighborhood of $S$.
Lemma 3.5 ([GS23, Remark 2.12]).
Any fanifold $\Phi$ admits a filtration
(3.1) $\Phi_{0}\subset\Phi_{1}\subset\cdots\subset\Phi_{n}=\Phi,\ n=\dim\Phi$
where $\Phi_{k}$ are subfanifolds defined as sufficiently small neighborhoods of $k$-skeleta $\operatorname{Sk}_{k}(\Phi)$, the closure of the subset of $k$-strata.
Proof.
Suppose that $\Phi_{k-1}$ has the desired filtration $\Phi_{0}\subset\Phi_{1}\subset\cdots\subset\Phi_{k-1}$, where $\Phi_{0}$ is the disjoint union $\bigsqcup_{P}\Sigma_{P}$ for all 00-strata $P\subset\Phi$ equipped with the canonically induced fanifold structure. The ideal boundary $\partial_{\infty}S$ might have some subset $\partial_{in}S$ which is in the direction to the interior of $\Phi$. Perform the gluing $\Phi_{k-1}\#_{(\Sigma_{S}\times\partial_{in}S)}(\Sigma_{S}\times S)$ which equals $\Sigma_{S}\times S$ when $\Phi_{k-1}$ is empty and equals $\Phi_{k-1}$ unless $S$ is an interior $k$-stratum, i.e., $\partial_{in}S=\partial S_{\circ}$ where $S_{\circ}=S\setminus\Phi_{k-1}$ is a manifold-with-boundary. Let
$\Phi_{k}=\Phi_{k-1}\#_{\bigsqcup_{S}(\Sigma_{S}\times\partial_{in}S)}\bigsqcup_{S}(\Sigma_{S}\times S)$
be the result of such gluings for all $k$-strata $S\subset\Phi$. Then $\Phi_{k}\subset\Phi$ is a subfanifold containing $\operatorname{Sk}_{k}(\Phi)$ since the products $\Sigma_{S}\times S$ are canonically fanifolds. ∎
3.2. Weinstein handle attachments
Let $W$ be a Weinstein domain with a smooth Legendrian $\mathcal{L}\subset\partial_{\infty}W$. Fix a standard neighborhood of $\mathcal{L}$ in $\partial_{\infty}W$
$\eta\colon\operatorname{Nbd}_{\partial_{\infty}W}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}=T^{*}\mathcal{L}\times\mathbb{R}$
which extends to a neighborhood in the Weinstein domain $W$
$\xi\colon\operatorname{Nbd}_{W}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}\cong T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0}),$
where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$. By [Eli, Proposition 2.9] one can modify the Weinstein structure near $\mathcal{L}$ so that the Liouville flow gets identified with the cotangent scaling on $T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$. We denote by $\widetilde{W}$ its conic completion. Then $\eta$ induces a neiborhood $\widetilde{\eta}\colon\operatorname{Nbd}_{\partial_{\infty}\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}$ which extends to a neighborhood $\widetilde{\xi}\colon\operatorname{Nbd}_{\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}$.
Remark 3.6.
From [GPS1, Lemma 2.13] it follows that $\widetilde{W}$ is a Weinstein sector.
Given Weinstein domains $W,W^{\prime}$ with smooth Legendrian embeddings $\partial_{\infty}W\hookleftarrow\mathcal{L}\hookrightarrow\partial_{\infty}W^{\prime}$, we write $W\#_{\mathcal{L}}W^{\prime}$ for the gluing
$W\#_{\eta^{-1}(T^{*}\mathcal{L}\times\{0\})\cong\eta^{\prime-1}(T^{*}\mathcal{L}\times\{0\})}W^{\prime}=\widetilde{W}\cup_{J^{1}\mathcal{L}}\widetilde{W^{\prime}}$
which yields a Weinstein sector with skeleton
$\operatorname{Core}(W\#_{\mathcal{L}}W^{\prime})=\operatorname{Core}(W,\mathcal{L})\cup_{\mathcal{L}}\operatorname{Core}(W^{\prime},\mathcal{L})$
where $\operatorname{Core}(W,\mathcal{L})$ is the relative skeleton of $(W,\lambda)$ associated with $\mathcal{L}$, i.e., the disjoint union $\operatorname{Core}(W)\sqcup\mathbb{R}\mathcal{L}\subset W$ of $\operatorname{Core}(W)$ and the saturation of $\mathcal{L}$ by the Liouville flow.
Definition 3.7 ([GS23, Definition 4.9, Lemma 4.10]).
Let $W$ be a Weinstein domain with a smooth Legendrian $\mathcal{L}\subset\partial_{\infty}W$. Then a Lagrangian $\mathbb{L}\subset W$ is biconic along $\mathcal{L}$ if and only if it is conic and
$\xi(\mathbb{L}\cap\operatorname{Nbd}_{W}(\mathcal{L}))\subset T^{*}\mathcal{L}\times\{0\}\times\mathbb{R}_{\leq 0}\subset J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}\cong T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$
for some $\xi\colon\operatorname{Nbd}_{W}(\mathcal{L})\hookrightarrow T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$, where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$.
Given biconic Lagrangians $\mathbb{L}\subset W,\mathbb{L}^{\prime}\subset W^{\prime}$ with matching ends
$\widetilde{\eta}(\mathbb{L}\cap\operatorname{Nbd}_{\partial_{\infty}W}(\mathcal{L}))=\widetilde{\eta}^{\prime}(\mathbb{L}^{\prime}\cap\operatorname{Nbd}_{\partial_{\infty}W^{\prime}}(\mathcal{L}))$
in $W\#_{\mathcal{L}}W^{\prime}$, we write $\mathbb{L}\#_{\mathcal{L}}\mathbb{L}^{\prime}$ for the gluing $\widetilde{\mathbb{L}}\cup_{\widetilde{\eta}(\mathbb{L}\cap\operatorname{Nbd}_{\partial_{\infty}W}(\mathcal{L}))}\widetilde{\mathbb{L}^{\prime}}$ in $W\#_{\mathcal{L}}W^{\prime}$, where $\widetilde{\mathbb{L}}\subset\widetilde{W},\widetilde{\mathbb{L}^{\prime}}\subset\widetilde{W^{\prime}}$ denote the saturations of $\mathbb{L},\mathbb{L}^{\prime}$ under the Liouville flows. Since any biconic subsets in $W,W^{\prime}$ remain conic in $W\#_{\mathcal{L}}W^{\prime}$, the gluing $\mathbb{L}\#_{\mathcal{L}}\mathbb{L}^{\prime}$ is a conic Lagrangian.
For a closed manifold $\widehat{M}$ and a compact manifold-with-boundary $S$, consider the Weinstein manifold-with-boundary $\widetilde{W^{\prime}}=T^{*}\widehat{M}\times T^{*}S$ with a smooth Legendrian $\mathcal{L}=\widehat{M}\times\partial S$ taken to be a subset of the zero section. Here, we equip $T^{*}S$ the tautological Liouville structure so that the Liouville vector field is the generator of fiberwise radial dilation. Then the Liouville flow on $\widetilde{W^{\prime}}$ near $\mathcal{L}$ is the cotangent scaling. We write $W^{\prime}$ for a Weinstein domain completing to $\widetilde{W^{\prime}}$. One can check that the above gluing procedure carries over, although $\mathcal{L}$ does not belong to $\partial_{\infty}W^{\prime}$.
Remark 3.8.
As explained in [GPS1, Example 2.3], when $S$ is noncompact $T^{*}S$ equipped with the tautological Liouville form is not even a Liouville manifold. However, if $S$ is the interior of a compact manifold-with-boundary, then $T^{*}S$ becomes a Liouville manifold after a suitable modification of the Liouville form near the boundary making it convex. We use the same symbol $T^{*}S$ to denote the result and regard it as a stopped Weinstein sector.
Definition 3.9 ([GS23, Definition 4.11]).
Let $W$ be a Weinstein domain with a smooth Legendrian embedding $\widehat{M}\times\partial S\hookrightarrow\partial_{\infty}W$.
(1) A handle attachment is the gluing $W\#_{\widehat{M}\times\partial S}W^{\prime}$ respecting the product structure.
(2) If $\mathbb{L}\subset W$ is a biconic subset along $\widehat{M}\times\partial S$ which locally factors as
$\eta(\mathbb{L}\cap\operatorname{Nbd}_{\partial_{\infty}W}(\widehat{M}\times\partial S))=\mathbb{L}_{S}\times\partial S\subset T^{*}\widehat{M}\times T^{*}\partial S$
for some fixed conic Lagrangian $\mathbb{L}_{S}$, then the extension of $\mathbb{L}$ through the handle is the gluing $\mathbb{L}\#_{\widehat{M}\times\partial S}(\mathbb{L}_{S}\times S)$ respecting the product structure.
Definition 3.10 ([FLTZ12, Section 3.1]).
Let $\Sigma\subset M_{\mathbb{R}}$ be a rational polyhedral fan and $\widehat{M}$ the real $n$-torus $\operatorname{Hom}(M,\mathbb{R}/\mathbb{Z})=M^{\vee}_{\mathbb{R}}/M^{\vee}$. The FLTZ Lagrangian is the union
$\mathbb{L}(\Sigma)=\bigcup_{\sigma\in\Sigma}\mathbb{L}_{\sigma}=\bigcup_{\sigma\in\Sigma}\sigma^{\perp}\times\sigma$
of conic Lagrangians $\sigma^{\perp}\times\sigma\subset\widehat{M}\times M_{\mathbb{R}}\cong T^{*}\widehat{M}$, where $\sigma^{\perp}$ is the real $(n-\dim\sigma)$-subtorus
$\{x\in\widehat{M}\ |\ \langle x,v\rangle=0\ \text{for all }v\in\sigma\}.$
Lemma 3.11 ([GS23, Lemma 4.16]).
Let $\Sigma\subset M_{\mathbb{R}}$ be a rational polyhedral fan. Then near the Legendrian $\partial_{\infty}\mathbb{L}_{\sigma}\subset\partial_{\infty}T^{*}\widehat{M}$ for each cone $\sigma\in\Sigma$ there are local coordinates
$\eta_{\sigma}\colon\operatorname{Nbd}(\partial_{\infty}\mathbb{L}_{\sigma})\hookrightarrow J^{1}\partial_{\infty}\mathbb{L}_{\sigma}=T^{*}\partial_{\infty}\mathbb{L}_{\sigma}\times\mathbb{R}$
with the following properties.
  • (1)
    The Lagranzian $\mathbb{L}_{\sigma}$ is biconic along $\partial_{\infty}\mathbb{L}_{\sigma}$.
  • (2)
    For any nonzero cones $\sigma,\tau\in\Sigma$ with $\sigma\subset\bar{\tau}$, we have
    (3.2) $\eta_{\sigma}(\partial_{\infty}\mathbb{L}_{\tau}\cap\operatorname{Nbd}(\partial_{\infty}\mathbb{L}_{\sigma}))=\mathbb{L}_{\tau/\sigma}\times\partial_{\infty}\sigma\times\{0\}\subset T^{*}\sigma^{\perp}\times T^{*}\partial_{\infty}\sigma\times\{0\}=T^{*}\partial_{\infty}\mathbb{L}_{\sigma}\times\{0\}.$
  • (3)
    The FLTZ Lagranzian $\mathbb{L}(\Sigma)$ is biconic along each $\partial_{\infty}\mathbb{L}_{\sigma}$.
  • (4)
    For each $\sigma\in\Sigma$ the local coordinates $\eta_{\sigma}$ define a Weinstein hypersuface
    $R_{\sigma}=\eta^{-1}_{\sigma}(T^{*}\partial_{\infty}\mathbb{L}_{\sigma}\times\{0\})\subset\partial_{\infty}T^{*}\widehat{M}$
    containing $\partial_{\infty}\mathbb{L}_{\sigma}$ as its skeleton and $R_{\tau}\cap\operatorname{Nbd}(\partial_{\infty}\mathbb{L}_{\sigma})\subset R_{\sigma}$ for any $\tau\in\Sigma$ with $\sigma\subset\bar{\tau}$.
Remark 3.12.
As explained above, we glue two Weinstein domains along the standard contact neighborhoods of their Weinstein hypersurfaces. Here, each $R_{\sigma}$ is a Weinstein thickening of $\partial_{\infty}\mathbb{L}_{\sigma}$. Lemma 3.11(4) guarantees compatibility of local Weinstein pairs $(T^{*}\widehat{M},R_{\sigma})$ with the union $\mathbb{L}(\Sigma)=\bigcup_{\sigma\in\Sigma}\mathbb{L}_{\sigma}$.
3.3. Construction
The stopped Weinstein sector $\widetilde{\mathbf{W}}(\Phi)$ is constructed inductively together with $\widetilde{\mathbb{L}}(\Phi)$ and $\pi$ along the filtration (3.1) of $\Phi$. When $k=0$, let $\widetilde{\mathbf{W}}(\Phi_{0})=\bigsqcup_{P}T^{*}\widehat{M}_{P},\ \widetilde{\mathbb{L}}(\Phi_{0})=\bigsqcup_{P}\mathbb{L}(\Sigma_{P})$ where each $T^{*}\widehat{M}_{P}$ is equipped with the canonical Liouville structure. We fix an identification $T^{*}\widehat{M}_{P}\cong\widehat{M}_{P}\times M_{P,\mathbb{R}}$ to regard each $\mathbb{L}(\Sigma_{P})$ as a conic Lagrangian submanifold of $T^{*}\widehat{M}_{P}$. Let $\pi_{0}\colon\widetilde{\mathbb{L}}(\Phi_{0})\to\Phi_{0}$ be the map induced by the projection to cotangent fibers. Then the triple $(\widetilde{\mathbf{W}}(\Phi_{0}),\widetilde{\mathbb{L}}(\Phi_{0}),\pi_{0})$ satisfies the conditions (1), …, (4). Define $\bar{\pi}_{0}$ as the composition $\operatorname{ret}_{0}\circ\tilde{\pi}_{0}$, where $\tilde{\pi}_{0}$ denotes the disjoint union of the projections to the cotangent fibers $T^{*}\widehat{M}_{P}\to M_{P,\mathbb{R}}$ and $\operatorname{ret}_{0}\colon\bigsqcup_{P}M_{P,\mathbb{R}}\to\bigsqcup_{P}\Sigma_{P}$ denotes the disjoint union of maps induced by retractions. See [Mor, Section 5.1] for details. The retractions cut off the part of the image of $\tilde{\pi}_{0}$ outside $\Phi$.
Suppose that we have constructed the triple $(\widetilde{\mathbf{W}}(\Phi_{k-1}),\widetilde{\mathbb{L}}(\Phi_{k-1}),\pi_{k-1})$ for the subfanifold $\Phi_{k-1}$. Let $\mathcal{L}_{k}$ be the disjoint union of
$\mathcal{L}^{k}_{S^{(l)}}=\pi^{-1}_{k-1}(S^{(l)})\cap\partial_{\infty}\widetilde{\mathbb{L}}(\Phi_{k-1})$
for all interior $l$-strata $S^{(l)}\subset\Phi,k\leq l$. There are smooth Legendrian embeddings
$\partial_{\infty}\mathbf{W}(\Phi_{k-1})\hookleftarrow\mathcal{L}^{k}_{S^{(l)}}\hookrightarrow\partial\bigsqcup_{S^{(k)}}(T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}).$
Due to the chart [GS23, Theorem 4.1(1)] for the previous step and Lemma 3.11(4), we may define $\widetilde{\mathbf{W}}(\Phi_{k})$ as the handle attachment
$\mathbf{W}(\Phi_{k-1})\#_{\mathcal{L}_{k}}\bigsqcup_{S^{(k)}}(T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}).$
There is a standard neighborhood $\eta_{k}\colon\operatorname{Nbd}_{\partial_{\infty}\mathbf{W}(\Phi_{k-1})}(\mathcal{L}_{k})\hookrightarrow J^{1}\mathcal{L}_{k}$ near $\mathcal{L}_{k}$ for which $\widetilde{\mathbb{L}}(\Phi_{k-1})$ is biconic along $\mathcal{L}_{k}$ and $\partial_{\infty}\widetilde{\mathbb{L}}(\Phi_{k-1})$ locally factors as
$\eta_{k}(\widetilde{\mathbb{L}}(\Phi_{k-1})\cap\operatorname{Nbd}_{\partial_{\infty}\mathbf{W}(\Phi_{k-1})}(\mathcal{L}_{k}))=\bigsqcup_{S^{(k)}}(\mathbb{L}(\Sigma_{S^{(k)}})\times\partial S^{(k)}_{\circ}).$
We define $\widetilde{\mathbb{L}}(\Phi_{k})$ as the extension through the disjoint union of the handles $T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}$
$\widetilde{\mathbb{L}}(\Phi_{k-1})\#_{\mathcal{L}_{k}}\bigsqcup_{S^{(k)}}(\mathbb{L}(\Sigma_{S^{(k)}})\times S^{(k)}_{\circ}).$
Let $\pi_{k}\colon\widetilde{\mathbb{L}}(\Phi_{k})\to\Phi_{k}$ be the map induced by $\pi_{k-1}$ and the projections from $T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}$ to the cotangent fiber direction in $T^{*}\widehat{M}_{S^{(k)}}$ and the base direction in $T^{*}S^{(k)}_{\circ}$. The triple $(\widetilde{\mathbf{W}}(\Phi_{k}),\widetilde{\mathbb{L}}(\Phi_{k}),\pi_{k})$ satisfies the conditions (1), …, (4).
Here, we explain how to construct $\bar{\pi}_{k}$ when attaching the handle $T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}$ to $\widetilde{\mathbf{W}}(\Phi_{k-1})$ for a single interior $k$-stratum $S^{(k)}\subset\Phi$ with $k\leq\dim\Phi$. Define $\tilde{\pi}_{k}$ as the map canonically induced by $\tilde{\pi}_{k-1},\tilde{\pi}_{0,S^{(k)}}$ and the projection $T^{*}S^{(k)}_{\circ}\to S^{(k)}_{\circ}$ to the base, where $\tilde{\pi}_{0,S^{(k)}}\colon T^{*}\widehat{M}_{S^{(k)}}\to M_{S^{(k)},\mathbb{R}}$ denotes the projection to the cotangent fibers. Here, we precompose the contraction
(3.3) $\operatorname{cont}_{k}\colon\widetilde{\mathbf{W}}(\Phi_{k})\to\widetilde{\mathbf{W}}(\Phi_{k})$
of the cylindrical ends along the negative Liouville flow to the part of $\partial_{\infty}\mathbf{W}(\Phi_{k-1})$ intersecting $\partial T^{*}S^{(k)}_{\circ}$. Let $\operatorname{ret}_{k}\colon M_{S^{(k)},\mathbb{R}}\to\Sigma_{S^{(k)}}$ be the map induced by a retraction defined similarly to $\operatorname{ret}_{0}$. Then $\bar{\pi}_{k}$ is the map canonically induced by $\bar{\pi}_{k-1},\bar{\pi}_{0,S^{(k)}}=\operatorname{ret}_{k}\circ\tilde{\pi}_{0,S^{(k)}}$ and the projection $T^{*}S^{(k)}_{\circ}\to S^{(k)}_{\circ}$ to the base.
4. Weinstein sectorial covers
In this section, we construct a Weinstein sectorial cover of $\widetilde{\mathbf{W}}(\Phi)$ by lifting a suitable cover of $\Phi$ along $\bar{\pi}$. The idea is similar to the construction for cotangent bundles. However, here we need extra care to make it compatible with the handle attachment process. Via sectorial descent, our cover gives rise to a constructible cosheaf $\bar{\pi}_{*}\operatorname{Fuk}_{*}$ of categories over $\Phi$ with respect to a certain natural topology. We show that $\bar{\pi}_{*}\operatorname{Fuk}_{*}$ is isomorphic to the $B$-side counterpart $\operatorname{Coh}_{!}\circ\mathbf{T}$.
4.1. Sectorial covers
Due to [GPS2], given a Liouville sector $X$, if it admits a Weinstein sectorial cover, then the wrapped Fukaya category $\mathcal{W}(X)$ can be computed by gluing that of pieces in the cover.
Definition 4.1 ([GPS2, Definition 12.2]).
Let $X$ be a Liouville manifold-with-boundary. A collection of cylindrical hypersurfaces $H_{1},\ldots,H_{n}\subset X$ is sectorial if their characteristic foliations $C_{1},\ldots,C_{n}$ are $\omega$-orthogonal over their intersections and there exist functions $I_{i}\colon\operatorname{Nbd}^{Z}_{X}(H_{i})\to\mathbb{R}$ linear near infinity satisfying
(4.1) $dI_{i}|_{C_{i}}\neq 0,\ dI_{i}|_{C_{j}}=0\ \text{for}\ i\neq j,\ \{I_{i},I_{j}\}=0.$
We also allow immersed hypersurfaces $H\to X$, with $I_{i}$ defined on $\operatorname{Nbd}^{Z}_{X}(H_{i})$ regarded as an immersed codimension zero submanifold of $X$, and the subscripts $i$ in the identities (4.1) indexing the “local branches” of $H$.
Definition 4.2 ([GPS2, Definition 12.14]).
A Liouville sector-with-sectorial-corners is a Liouville manifold-with-corners whose boundary, regarded as an immersed hypersurface, is sectorial.
Definition 4.3 ([GPS2, Definition 12.19, Remark 12.20]).
Let $X$ be a Liouville manifold-with-boundary. Suppose that $X$ admits a cover $X=\bigcup^{n}_{i=1}X_{i}$ by Liouville sectors-with-sectorial-corners $X_{i}\subset X$, with precisely two faces $\partial^{1}X_{i}=X_{i}\cap\partial X$ and the point set topological boundary $\partial^{2}X_{i}$ of $X_{i}$ meeting along the corner locus $\partial X\cap\partial^{2}X_{i}=\partial^{1}X_{i}\cap\partial^{2}X_{i}$. Such a cover is sectorial if the collection of boundaries $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ is sectorial. We also allow $X_{i}$ without corners, requiring $\partial^{2}X_{i}$ to be disjoint from $\partial X$, $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ to be sectorial and $\bigcup^{n}_{i=1}X_{i}\hookrightarrow X$ to be a trivial inclusion.
For any sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$, stratify $X$ by strata
(4.2) $X_{J,K,L}=\bigcup_{j\in J}X_{j}\cap\bigcap_{k\in K}\partial X_{k}\ \setminus\bigcap_{l\in L}X_{l}$
ranging over $J\sqcup K\sqcup L=\{1,\ldots,n\}$. The closure of each $X_{J,K,L}$ is a submanifold-with-corners, whose symplectic reduction is a Liouville sector-with-sectorial-corners.
Definition 4.4 ([GPS2, Definition 12.15]).
A sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ is Weinstein if the convex completions of all of the symplectic reductions of strata (4.2) are Weinstein up to deformation.
Theorem 4.5 ([GPS2, Theorem 1.35]).
For any Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ of a Weinstein sector $X$, the induced functor
(4.3) $\displaystyle\operatorname{colim}_{\emptyset\neq I\subset\{1,\ldots,n\}}\mathcal{W}(\bigcap_{i\in I}X_{i})\to\mathcal{W}(X)$
is a pretriangulated equivalence. Moreover, the same holds for any mostly Legendrian stop $\mathfrak{r}\subset(\partial_{\infty}X)^{\circ}$ which is disjoint from each $\partial X_{i}$ and from a neighborhood of $\partial X$.
4.2. Alternative description of $\widetilde{\mathbf{W}}(\Phi)$ and $\bar{\pi}$
Recall that $\widetilde{\mathbf{W}}(\Phi)$ is a Weinstein manifold-with-boundary obtained by Weinstein handle attachments, special cases of gluing of Weinstein pairs. The boundary comes from conic completion of the boundaries of the contact surroundings of $T^{*}\mathcal{L}_{k}$ in $\partial_{\infty}{\mathbf{W}}(\Phi_{k-1})$. Below, preserving a sufficiently small neighborhood of $\widetilde{\mathbb{L}}(\Phi)$, we will modify the construction in order to obtain a Weinstein manifold without boundary. See [Eli, Figure 3.1] to get the picture of our modification. As in the original construction, we proceed by induction on $k$.
Whenever we attach a handle $T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}$ to $\widetilde{\mathbf{W}}(\Phi_{k-1})$, deform $\partial_{\infty}\widetilde{\mathbf{W}}(\Phi_{k-1})$ so that the complement in $\partial_{\infty}\widetilde{\mathbf{W}}(\Phi_{k-1})$ of the contact surrounding $T^{*}\mathbb{L}(\Sigma_{S^{(k)}})\times\partial S^{(k)}_{\circ}$ gets connected smoothly to $T^{*}\widehat{M}_{S^{(k)}}\times\partial_{\infty}T^{*}S^{(k)}_{\circ}$. Here, we regard $T^{*}S^{(k)}_{\circ}$ as a Liouville sector equipped with the tautological Liouville form. Now, conic completion yields a Weinstein manifold without boundary. We use the same symbol $\widetilde{\mathbf{W}}(\Phi)$ to denote the result, as our modification preserves a sufficiently small neighborhood of $\widetilde{\mathbb{L}}(\Phi)$. One can also modify the map $\bar{\pi}$ in a straightforward way. Namely, when precomposing (3.3), we contract the cylindrical ends of $T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ}$ to $T^{*}\widehat{M}_{S^{(k)}}\times\partial_{\infty}T^{*}S^{(k)}_{\circ}$.
4.3. Construction
We construct a Weinstein sectorial cover of the modification $\widetilde{\mathbf{W}}(\Phi)$ as the inverse image under $\bar{\pi}$ of a certain cover of $\Phi$. For expository reason, we will assume that $\Phi$ is closed. Since each stratum is closed, one can take its barycenter. Connect by an edge inside $\Phi$ the barycenter on each top dimensional stratum with that on $1$-dimensional lower strata adjacent to it. Repeat the process inductively until reaching the barycenters of $1$-strata. There might be some lower dimensional closed strata which are not adjacent to higher dimensional strata. Beginning the process also from such strata, we obtain a partition of $\Phi$ defined by the additional edges connecting the barycenters and the barycenters on those $1$-strata which are not adjacent to higher dimensional strata. The partition divides $\Phi$ into some finite number of connected components.
For each 00-stratum $P$, there is a unique connected component containing $P$. Take a suitable slight enlargement of the rounding the corners of its closure. We denote by $\widetilde{\mathbf{W}}(P)$ and $\partial^{2}\widetilde{\mathbf{W}}(P)$ the inverse image under $\bar{\pi}$ respectively of the slight enlargement and of its boundary. Perturbing these boundaries if necessary, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ are disjoint or intersect transversely for any pair $P,P^{\prime}$ of 00-strata. Perturb each edge connecting the barycenter of a top dimensional stratum $S^{(n)}$ to that of a $1$-dimensional lower stratum $S^{(n-1)}$ so that it becomes locally in $\Phi_{n-1}$ perpendicular to $S^{(n-1)}$. More precisely, the edge becomes parallel to the normal direction in the handle $\Sigma_{S^{(n-1)}}\times S^{(n-1)}\subset\Phi_{n-1}\setminus\Phi_{n-2}$ to $S^{(n-1)}$. This is always possible since $\Phi$ is smoothly normally conical. Repeat the process inductively until reaching the barycenters of $1$-strata in $\Phi_{1}$. Perform the same perturbation to each edge starting from the barycenter of a lower dimensional closed stratum which is not adjacent to higher dimensional strata. We use the same symbols $\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P)$ to denote the results.
Remark 4.6.
Our construction works even when $\Phi$ is not closed. Then for each stratum we will take the barycenter of its closure, which makes sense due to the condition $(iii)$. Moreover, for instance, we will have connected components not adjacent to vertices, which only causes irrelevant complication in labeling.
Lemma 4.7.
Each $\widetilde{\mathbf{W}}(P)$ is a Weinstein sector.
Proof.
We adapt the idea in [GPS2, Example 1.33]. Recall that we are assuming $\Phi$ to be closed. For simplicity, we will further assume that there is no lower dimensional strata which are not adjacent to higher dimensional strata. Since by construction $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ is away from $P$, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is disjoint from $T^{*}\widehat{M}_{P}\subset\widetilde{\mathbf{W}}(P)\cap\widetilde{\mathbf{W}}(\Phi_{0})$. Let $I$ be a $1$-stratum which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ at a point $p$. Due to the above perturbation, locally in $\Phi_{1}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $I$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{I}$ and the fiber direction of $T^{*}I_{\circ}$, over the neighborhood of $I$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{I}$ and the fiber $T^{*}_{p}I_{\circ}$. Inductively, let $S^{(k)}$ be a $k$-stratum adjacent to a $(k-1)$-stratum $S^{(k-1)}$ which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$. Due to the above perturbation, locally in $\Phi_{k}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $S^{(k)}$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{S^{(k)}}$ and the fiber direction of $T^{*}S^{(k)}_{\circ}$, over the neighborhood of $S^{(k)}$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{S^{(k)}}$ and the restriction $T^{*}S^{(k)}_{\circ}|_{\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}}$. By construction locally in $\Phi_{(k-1)}$ the intersection $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}$ becomes perpendicular to $S^{(k-1)}$, the handle attachment $\mathbf{W}(\Phi_{k-1})\#_{\mathcal{L}_{k-1}}(T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ})$ glues this product to the gluing of the products in the previous steps.
Rewinding the process, one sees that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is obtained by extending via the handle attachments the union of the inverse images under $\bar{\pi}$ of hypersurfaces in the intersection of $\Phi_{n}\setminus\Phi_{n-1}$ and the connected component containing $P$. We claim that this hypersurface $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial. Since the newly attached pieces are cylindrical, so is $\partial^{2}\widetilde{\mathbf{W}}(P)$. It remains to show the existence of an $\alpha$-defining function on $\partial^{2}\widetilde{\mathbf{W}}(P)$. By [GPS2, Construction 12.17] we may assume $\partial^{2}\widetilde{\mathbf{W}}(P)$ to be one of the hypersurfaces-with-corners dividing $\widetilde{\mathbf{W}}(\Phi)$ into the original connected components. Let $I$ be a $1$-stratum adjacent to $P$ and $H_{P,I}$ any of such hypersurfaces intersecting $I$. Fix the standard coordinates $(x,y)$ on the symplectic manifold $T^{*}I_{\circ}$. Unwinding the filtration of $\Phi$ and the construction of $\widetilde{\mathbf{W}}(\Phi)$, one sees that $H_{P,I}$ is given by $\{x=0\}$ up to perturbation. We denote by $y_{P,I}$ the function on a neighborhood of $H_{P,I}$ which is canonically induced by $y\colon T^{*}I_{\circ}\to\mathbb{R}$. Clearly, $y_{P,I}$ is linear near infinity. The Hamiltonian vector field $X_{y_{P,I}}$ is outward pointing along $H_{P,I}$, as it is given by a lift of $\partial_{x}$ on $T^{*}I_{\circ}$. Hence $H_{P,I}$ is sectorial. Thus the Weinstein manifold-with-boundary $\widetilde{\mathbf{W}}(P)$ satisfies the condition in Definition 2.1 to be a Liouville sector. ∎
Remark 4.8.
Suppose that there is a lower dimensional closed stratum which is not adjacent to higher dimensional strata. Let $S^{(k)}$ be such a stratum. Then for each 00-stratum $P$ adjacent to it, $\partial^{2}\widetilde{\mathbf{W}}(P)$ gets contribution from the extensions via the handle attachments of the inverse images under $\bar{\pi}$ of hypersurfaces in $S^{(k)}$.
Remark 4.9.
Suppose that $\Phi$ is not closed and, for instance, there is a half-open $1$-stratum $I$ adjacent to $P$. Recall that in this case the handle attachment $\mathbf{W}(\Phi_{1})\#_{\mathcal{L}_{1}}(T^{*}\widehat{M}_{I}\times T^{*}I_{\circ})$ is not precisely a Liouville manifold-with-boundary, as one needs to modify the canonical Liouville form of $T^{*}I_{\circ}$ near $\partial T^{*}I_{\circ}$ making it convex. This modification, which is not necessary if $I$ were closed, yields a stop on the newly formed ideal boundary of the handle attachment.
Lemma 4.10.
The union $\bigcup_{P}\widetilde{\mathbf{W}}(P)$ gives a Weinstein sectorial cover of the modification $\widetilde{\mathbf{W}}(\Phi)$.
Proof.
We adapt the idea in [GPS2, Example 1.34]. Again, assume that $\Phi$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata. Finiteness of the cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ follows from the condition $(i)$. By construction each $\widetilde{\mathbf{W}}(P)$ has no corners and $\partial\widetilde{\mathbf{W}}(\Phi)$ is empty. Moreover, $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ intersect transversely for any distinct pair $P,P^{\prime}$ of 00-strata unless they are disjoint. By Lemma 4.7 each $\widetilde{\mathbf{W}}(P)$ is a Weinstein sector. Hence the collection of the hypersurfaces $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial if there exist functions
$\tilde{y}_{P}\colon\operatorname{Nbd}^{Z}_{\widetilde{\mathbf{W}}(\Phi)}(\partial^{2}\widetilde{\mathbf{W}}(P))\to\mathbb{R}$
linear near infinity satisfying
$d\tilde{y}_{P}|_{C_{P^{\prime}}}=0\ \text{for}\ P\neq P^{\prime},\ \{\tilde{y}_{P},\tilde{y}_{P^{\prime}}\}=0$
where $C_{P}$ denote the characteristic foliations of $\partial^{2}\widetilde{\mathbf{W}}(P)$.
Consider the functions $y_{P,I}$ in the proof of Lemma 4.7. Pushing forward $X_{y_{P,I}}$ by $\bar{\pi}$, we obtain a collection of vector fields on neighborhoods of $\bar{\pi}(H_{P,I})$. Here, by a vector field on a subset of $\Phi$ we mean the gluing of vector fields on submanifolds of $\Phi_{k}\setminus\Phi_{k-1}$ for $k=1,\ldots,n$. Take their deformations around the corners to concatenate $\bar{\pi}_{*}(X_{P,I})$ so that translation of the result gives a vector field $\tilde{X}_{P}$ on a neighborhood $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$, which is outward pointing along $\partial^{2}\widetilde{\mathbf{W}}(P)$. Recall that $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P)),\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P^{\prime}))$ intersect transversely for any distinct pair $P,P^{\prime}$ of 00-strata unless they are disjoint. Since $\Phi$ has only finitely many 00-strata, we may deform further $\tilde{X}_{P}$ simultaneously so that they in addition become tangent to $\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ on the intersections. We use the same symbol $\tilde{X}_{P}$ to denote these deformations. We define $\tilde{y}_{P}$ as the Hamiltonian lifts of $\tilde{X}_{P}$, which make sense due to the construction of $\widetilde{\mathbf{W}}(\Phi)$ and our definition of a vector field on a subset of $\Phi$.
It remains to show that the sectorial cover is Weinstein. All strata
$\widetilde{\mathbf{W}}(\Phi)_{J,K,L}=\bigcup_{P\in J}\widetilde{\mathbf{W}}(P)\cap\bigcap_{P^{\prime}\in K}\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})\setminus\bigcap_{P^{\prime\prime}\in L}\widetilde{\mathbf{W}}(P^{\prime\prime})$
are the inverse images under $\bar{\pi}$ of
$\bar{\pi}(\widetilde{\mathbf{W}}(\Phi))_{J,K,L}=\bigcup_{P\in J}\bar{\pi}(\widetilde{\mathbf{W}}(P))\cap\bigcap_{P^{\prime}\in K}\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P^{\prime}))\setminus\bigcap_{P^{\prime\prime}\in L}\bar{\pi}(\widetilde{\mathbf{W}}(P^{\prime\prime}))$
Recall that locally in $\Phi_{k}$ the images $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ are perpendicular to $S^{(k)}$. Hence over $\Phi_{k}\setminus\Phi_{k-1}$ the convex completions of the symplectic reductions of $\widetilde{\mathbf{W}}(\Phi)_{J,K,L}$ are given by the union of products of $T^{*}\widehat{M}_{S^{(k)}}$ and the restrictions $T^{*}S^{(k)}_{\circ}|_{\bar{\pi}(\widetilde{\mathbf{W}}(\Phi))_{J,K,L}\cap S^{(k)}}$. In particular, over $\Phi_{n}\setminus\Phi_{n-1}$ the convex completions of the symplectic reductions of $\widetilde{\mathbf{W}}(\Phi)_{J,K,L}$ are simply given by the cotangent bundles of $\bar{\pi}(\widetilde{\mathbf{W}}(\Phi))_{J,K,L}$. These unions for $k=1,\ldots,n$ are Weinstein up to deformation and glued to yield the convex completion of the symplectic reduction of $\widetilde{\mathbf{W}}(\Phi)_{J,K,L}$. ∎
Remark 4.11.
The construction does not work without modification of $\widetilde{\mathbf{W}}(\Phi)$, as in general $\partial\widetilde{\mathbf{W}}(\Phi)$ is nonempty and intersects $\partial^{2}\widetilde{\mathbf{W}}(P)$.
4.4. Sectorial descent and HMS over fanifolds
Our Weinstein sectorial cover is compatible with HMS established by Gammage–Shende in [GS23]. On the $A$-side, there are three isomorphic constructible cosheaves of categories over $\Phi$. First, consider the functor
$\operatorname{fsh}_{*}\colon(\operatorname{Fan}^{\twoheadrightarrow})^{op}\to^{**}\operatorname{DG},\ \Sigma\mapsto\operatorname{Sh}_{\mathbb{L}(\Sigma)}(T_{\Sigma})$
which is obtained from $\operatorname{fsh}\colon\operatorname{Fan}^{\twoheadrightarrow}\to^{*}\operatorname{DG}^{*}$ in [GS23, Section 4.5] by taking adjoints of the images of the morphisms. Here, ${}^{**}\operatorname{DG}$ denotes the category of cocomplete dg categories and functors which preserves colimits and compact objects, while ${}^{*}\operatorname{DG}^{*}$ denotes the category of cocomplete dg categories and functors which preserves limits and colimits. Note that we have ${}^{*}\operatorname{DG}^{*}=(^{**}\operatorname{DG})^{op}$. We use the same symbol $\operatorname{fsh}_{*}$ to denote the composition $\operatorname{Exit}(\Phi)^{op}\to^{**}\operatorname{DG}$ with the functor $\operatorname{Exit}(\Phi)^{op}\to(\operatorname{Fan}^{\twoheadrightarrow})^{op}$ equipped with $\Phi$.
Second, consider the functor
$\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)*}\colon\operatorname{Exit}(\Phi)^{op}\to^{**}\operatorname{DG}$
which is obtained from $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)}\colon\operatorname{Exit}(\Phi)\to^{*}\operatorname{DG}^{*}$ in [GS23, Section 4.5] by taking adjoints of the images of the morphisms. By definition both $\operatorname{fsh}_{*}$ and $\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)*}$ are constructible cosheaves of categories over $\Phi$.
Lemma 4.12 (cf. [GS23, Proposition 4.34]).
There is an isomorphism
$\operatorname{fsh}_{*}\cong\pi_{*}\operatorname{\mu sh}_{\widetilde{\mathbb{L}}(\Phi)*}.$
Proof.
The claim immediately follows from [GS23, Proposition 4.34]. ∎
Third, let $\operatorname{UI}(\Phi)$ be the category of unions of intersections of $\widetilde{\mathbf{W}}(P)$ in the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ from Lemma 4.10. Morphisms are given by inclusions of Weinstein sectors. Introduce to $\operatorname{UI}(\Phi)$ a topology generated by maximal proper Weinstein subsectors $\widetilde{\mathbf{W}}(P)$ of $\widetilde{\mathbf{W}}(\Phi)$ in the cover. Consider the functor
$\operatorname{Fuk}_{*}\colon\operatorname{UI}(\Phi)\to^{**}\operatorname{DG}$
which sends each Weinstein sector to the category of modules over the wrapped Fukaya category. The images of morphisms are induced by the pushforward functors from [GPS1] along the inclusions of Weinstein sectors. Due to the covariant functoriality and sectorial descent of wrapped Fukaya categories, $\operatorname{Fuk}_{*}$ gives a cosheaf of categories over $\operatorname{UI}(\Phi)$. Note that $\operatorname{Fuk}_{*}$ preserves compact objects as it is defined before taking module categories. Since $\bar{\pi}$ restricts to $\pi$ and $\widetilde{\mathbf{W}}(\Phi)=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a lift of an open cover of $\widetilde{\mathbb{L}}(\Phi)$, the pushforward
$\bar{\pi}_{*}\operatorname{Fuk}_{*}\colon\operatorname{Exit}(\Phi)^{op}\to^{**}\operatorname{DG}$
gives a constructible cosheaf of categories over $\Phi$.
Lemma 4.13.
There is an isomorphism
$\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\bar{\pi}_{*}\operatorname{Fuk}_{*}.$
Proof.
Over a sufficiently small open neighborhood of each intersection of $\bar{\pi}(\widetilde{\mathbf{W}}(P))$, the inverse image under $\bar{\pi}$, regarded as a stopped Weinstein sector as in Remark 3.8, has the inverse image under $\pi$ as its relative skeleton. Hence by [GPS3, Theorem 1.4] the sections over such open neighborhoods are canonically equivalent. Moreover, the corestriction functors between sections of the left and the pushforward functors between sections of the right intertwine these canonical equivalences. ∎
On the $B$-side, the composition
$\operatorname{Coh}_{!}\circ\mathbf{T}\colon\operatorname{Exit}(\Phi)^{op}\to^{**}\operatorname{DG},\ S^{(k)}\mapsto\operatorname{Coh}(\mathbf{T}(S^{(k)}))$
of $\operatorname{Coh}_{!}$ with $\mathbf{T}$ from [GS23, Section 3] gives a constructible cosheaves of categories over $\Phi$. Recall that $\mathbf{T}(S^{(k)})$ is the toric variety associated with the fan $\Sigma_{S^{(k)}}\subset M_{S^{(k)},\mathbb{R}}$, where $(\Sigma_{S^{(k)}},M_{S^{(k)}})$ is the image of $S^{(k)}$ under the functor $\operatorname{Exit}(\Phi)^{op}\to(\operatorname{Fan}^{\twoheadrightarrow})^{op}$ equipped with $\Phi$. Recall also that the values of $\operatorname{Coh}_{!}\circ\mathbf{T}$ are the module categories over the dg categories of coherent sheaves on toric varieties and the images of morphisms are the pushforwards along closed immersions.
Lemma 4.14 ([GS23, Proposition 3.10]).
The colimit
$\mathbf{T}(\Phi)=\operatorname{colim}_{\operatorname{Exit}(\Phi)^{op}}\mathbf{T}(S^{(k)})$
exists as an algebraic space.
Lemma 4.15 ([GS23, Proposition 3.14]).
The canonical functor
$\operatorname{Coh}_{!}\circ\mathbf{T}(\Phi)=\operatorname{colim}_{\operatorname{Exit}(\Phi)^{op}}\operatorname{Coh}\mathbf{T}(S^{(k)})\to\operatorname{Coh}\left(\operatorname{colim}_{\operatorname{Exit}(\Phi)^{op}}\mathbf{T}(S^{(k)})\right)=\operatorname{Coh}(\mathbf{T}(\Phi))$
is an equivalence.
Now, it is straightforward to describe HMS in terms of cosheaves of categories over $\Phi$.
Lemma 4.16 (cf. [GS23, Theorem 5.3]).
There is an isomorphism
$\pi_{*}\operatorname{\mu sh}^{op}_{\widetilde{\mathbb{L}}(\Phi)*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}.$
Proof.
The claim follows immediately from [GS23, Theorem 5.3]. ∎
Theorem 4.17 (cf. [GS23, Corollary 5.8]).
There is an isomorphism
$\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}.$
Proof.
Concatenate Lemma 4.13 and Lemma 4.16. ∎
Taking the global section, we obtain
Corollary 4.18 (cf. [GS23, Theorem 5.4]).
There is an equivalence
(4.4) $\operatorname{Fuk}(\widetilde{\mathbf{W}}(\Phi))\simeq\operatorname{Coh}(\mathbf{T}(\Phi)).$
5. HMS for very affine hypersurfaces
In this section, we revisit HMS for very affine hypersurfaces established by Gammage–Shende. Specifically, we describe an aspect of the mirror pair as the associated spaces with a fanifold. Then we apply our results in the previous section, giving an affirmative answer to the conjecture raised in [GS22, Section 2.4]. As a consequence, we promote HMS for very affine hypersurfaces to an isomorphism of constructible cosheaves of categories over the fanifold.
5.1. Mirror pair for very affine hypersurfaces
To a very affine hypersuface $H\subset M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}\cong(\mathbb{C}^{*})^{d+1}$ satisfying certain assumptions on the defining Laurent polynomial, the toric boundary divisor $\partial\mathbf{T}_{\Sigma}$ gives its mirror [GS22]. Here, $\mathbf{T}_{\Sigma}$ is the associated toric stack with a smooth quasiprojective stacky fan $\Sigma\subset M_{\mathbb{R}}\cong\mathbb{R}^{d+1}$. Recall that $\Sigma$ is determined by the defining polynomial of $H$. Under the assumptions, its stacky primitives span a convex lattice polytope $\Delta^{\vee}$ containing the origin and $\Sigma$ defines an adapted star-shaped triangulation $\mathcal{T}$ of $\Delta^{\vee}$. Below, we review the proof of the following statement.
Theorem 5.1 ([GS22, Theorem 1.0.1]).
Assume that the Newton polytope $\Delta^{\vee}$ contains the origin and its triangulation $\mathcal{T}$ is star-shaped. Then there is an equivalence
(5.1) $\operatorname{Fuk}(H)\simeq\operatorname{Coh}(\partial\mathbf{T}_{\Sigma}).$
Due to the assumptions, one can easily construct a global skeleton $\operatorname{Core}(H)$ of $H$. Recall that we restrict to $H$ the canonical Weinstein structure on $M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}$. In the sequel, for simplicity we will assume that $\Sigma$ is an ordinary simplicial fan. When $\Sigma$ is a smooth quasiprojective stacky fan, we will replace the $A$-model with its finite cover and the $B$-model with the associated toric stack. The skeleton $\operatorname{Core}(H)$ can be obtained by gluing skeleta of tailored pants $\tilde{P}_{d}$ computed by Nadler [Nad], after transported to the ones in the pants decomposition of $H$. As explained in [Nad], we may identify the skeleton $\operatorname{Core}(\tilde{P}_{d})$ with the ideal boundary of the FLTZ skeleton associated with the toric variety $\mathbb{A}^{d+1}$. Hence we may identify $\operatorname{Core}(H)$ with $\partial_{\infty}\mathbb{L}(\Sigma)$ for the FLTZ skeleton $\mathbb{L}(\Sigma)$ associated with $\mathbf{T}_{\Sigma}$.
Lemma 5.2 ([GS22, Corollary 4.3.2]).
Assume that the Newton polytope $\Delta^{\vee}$ contains the origin and its triangulation $\mathcal{T}$ is star-shaped. Then $\partial_{\infty}\mathbb{L}(\Sigma)$ has an open cover by the subsets $\mathbb{L}(\Sigma/\sigma)\times\partial_{\infty}\sigma$ anti-indexed by the poset of nonzero cones in $\Sigma$.
Proof.
Let $\sigma\subset\sigma^{\prime}\subset\tau$ be nonzero cones in $\Sigma$. Then $\tau/\langle\sigma\rangle,\tau/\langle\sigma^{\prime}\rangle$ give cones in the normal fans $\Sigma/\sigma,\Sigma/\sigma^{\prime}$ with the isomorphisms
$(\tau/\langle\sigma\rangle)^{\perp}\times\tau/\langle\sigma\rangle\times\sigma\cong(\tau)^{\perp}\times\tau\cong(\tau/\langle\sigma^{\prime}\rangle)^{\perp}\times\tau/\langle\sigma^{\prime}\rangle\times\sigma^{\prime}.$
Note that subtori $(\tau/\langle\sigma\rangle)^{\perp},(\tau/\langle\sigma^{\prime}\rangle)^{\perp}$ of respective tori $\widehat{M}/\langle\sigma\rangle,\widehat{M}/\langle\sigma^{\prime}\rangle$ coincide with $\tau^{\perp}\subset\widehat{M}$. Hence we obtain an open inclusion
$\mathbb{L}(\Sigma/\sigma^{\prime})\times\sigma^{\prime}\hookrightarrow\mathbb{L}(\Sigma/\sigma)\times\sigma.$
Now, the claim follows from $\partial_{\infty}(\mathbb{L}(\Sigma/\sigma)\times\sigma)\cong\mathbb{L}(\Sigma/\sigma)\times\partial_{\infty}\sigma$. ∎
To each $\sigma\in\Sigma$, let $\overline{O(\sigma)}$ be the closure of the corresponding toric orbit $O(\sigma)$. It is the toric variety associated with $\Sigma/\sigma$. Here, we take its reduced induced scheme structure. If $\sigma\subset\tau$, then there is a closed immersion $\overline{O(\tau)}\hookrightarrow\overline{O(\sigma)}$ induced by the quotient $\Sigma/\sigma\to\Sigma/\tau$. Also, we have $\overline{O(\sigma)}\cap\overline{O(\sigma^{\prime})}=\overline{O(\sigma\wedge\sigma^{\prime})}$ where $\sigma\wedge\sigma^{\prime}$ denotes the smallest cone in $\Sigma$ containing both $\sigma$ and $\sigma^{\prime}$ if such a cone exists, otherwise $\emptyset$. Recall that by definition $\partial\mathbf{T}_{\Sigma}$ is the union of the nontrivial toric orbit closures. Scheme-theoretically, this amounts to
Lemma 5.3 ([GS22, Lemma 3.4.1]).
The colimit
$\partial\mathbf{T}_{\Sigma}=\operatorname{colim}_{0\neq\sigma\in\Sigma}\overline{O(\sigma)}$
exists as an algebraic space.
From the above lemma it follows
$\operatorname{Coh}(\partial\mathbf{T}_{\Sigma})=\operatorname{Coh}(\operatorname{colim}_{0\neq\sigma\in\Sigma}\overline{O(\sigma)})=\operatorname{colim}_{0\neq\sigma\in\Sigma}\operatorname{Coh}(\overline{O(\sigma)}).$
Now, one can easily obtain (5.1). Indeed, by [GPS3, Theorem 1.4] and Lemma 5.2 we have
$\operatorname{Fuk}(H)=\operatorname{\mu sh}_{\mathbb{L}(\Sigma)}(\mathbb{L}(\Sigma))^{op}=\operatorname{colim}_{0\neq\sigma\in\Sigma}\operatorname{\mu sh}_{\mathbb{L}(\Sigma)}(\sigma^{\perp}\times\partial_{\infty}\sigma)^{op}.$
As explained in the proof of [GS22, Theorem 7.4.1], we have
$\operatorname{colim}_{0\neq\sigma\in\Sigma}\operatorname{\mu sh}_{\mathbb{L}(\Sigma)}(\sigma^{\perp}\times\partial_{\infty}\sigma)^{op}=\operatorname{colim}_{0\neq\sigma\in\Sigma}\operatorname{Coh}(\overline{O(\sigma)}).$
5.2. Mirror pair over a fanifold
Up to tailoring $H$, the mirror pair $(H,\partial\mathbf{T}_{\Sigma})$ can be obtained as a special case of mirror pairs $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ over $\Phi$.
Example 5.4 ([GS23, Example 2.11, 4.23]).
Consider the standard embedding
$S^{d}=\{\Sigma^{d+1}_{i=1}x^{2}_{i}=1\}\subset\mathbb{R}^{d+1}$
of the $d$-sphere. Then $\Phi_{\Sigma}=\Sigma\cap S^{d}\subset S^{d}$ carries a canonical fanifold structure. For instance, any vertex $P$ is the intersection of a ray $\rho_{P}\in\Sigma(1)$ and $S^{d}$. To it one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$. Clearly, $\Phi_{\Sigma}$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata.
Lemma 5.5.
There are open neighborhoods of $\operatorname{Core}(H)$ in $H$ and $\widetilde{\mathbb{L}}(\Phi_{\Sigma})$ in $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ which are isomorphic as a Weinstein manifold.
Proof.
For each top dimensional cone $\sigma\in\Sigma_{\max}$, consider the subfanifold $\Phi_{\sigma}=\bar{\sigma}\cap S^{d}$ of $\Phi_{\Sigma}$. Unwinding the proof of [Nad, Theorem 5.13], one sees that the associated Weinstein sector $\widetilde{\mathbf{W}}(\Phi_{\sigma})$ can be identified with the open neighborhood $U_{d}$ there of $\operatorname{Core}(\tilde{P}_{d})$ in $\tilde{P}_{d}$, after we canonically transform $\tilde{P}_{d}$ to the pants whose tropicalization is dual to $\sigma$. Hence, up to suitable transformation, we may regard $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ as the union of $U_{d}$ and $\widetilde{\mathbb{L}}(\Phi_{\Sigma})$ as the union of $\operatorname{Core}(\tilde{P}_{d})$. Then, near their skeleta, $H$ and $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ are symplectomorphic by the same argument as in [Nad, Theorem 5.13]. Since over $\Phi_{\Sigma}$ the Weinstein handle attachments can be carried out in the ambient cotangent bundle $M^{\vee}_{\mathbb{C}}\cong T^{*}T^{d+1}$, the Weinstein structure on $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ coincides with the restriction of the canonical one on $M^{\vee}_{\mathbb{C}}$. ∎
Lemma 5.6.
The algebraic space $\mathbf{T}(\Phi_{\Sigma})$ coincides with $\partial\mathbf{T}_{\Sigma}$.
Proof.
By Lemma 4.14 we have $\mathbf{T}(\Phi_{\Sigma})=\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})$. As one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$ to each vertex $P\in\Phi_{\Sigma}$, the toric variety $\mathbf{T}(P)$ coincides with $\overline{O(\rho_{P})}$. Let $\sigma_{S^{(k)}}\in\Sigma_{P}$ be the cone corresponding to an exit path in $\Phi_{\Sigma}$ from $P$ to $S^{(k)}$. Then by definition of fanifolds $\mathbf{T}(S^{(k)})$ is the toric variety associated with the fan $\Sigma_{P}/\sigma_{S^{(k)}}\subset M_{\mathbb{R}}/\langle\sigma_{S^{(k)}}\rangle$, which in turn is the toric orbit closure corresponding to the inverse image $\tilde{\sigma}_{S^{(k)}}$ of $\sigma_{S^{(k)}}$ under the quotient $\Sigma\to\Sigma_{P}$. When $S^{(k)}$ runs through the strata of $\Phi_{\Sigma}$, the inverse images $\tilde{\sigma}_{S^{(k)}}$ range over nonzero cones in $\Sigma$. Thus we have
$\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})=\operatorname{colim}_{0\neq\tilde{\sigma}_{S^{(k)}}\in\Sigma}\overline{O(\tilde{\sigma}_{S^{(k)}})}.$
Now, the claim follows from Lemma 5.3. ∎
5.3. Sectorial descent and HMS for very affine hypersufaces
By Lemma 5.5 and Lemma 5.6 the equivalences (5.1) and (4.4) form a commutative diagram
[…]
where the lower equivalence is the global section of the isomorphism $\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$ of constructible cosheaves of categories over $\Phi_{\Sigma}$ from Theorem 4.17. Finally, we give an affirmative answer to the conjecture raised in [GS22, Section 2.4].
Theorem 5.7.
There is a Weinstein sectorial cover of $H$ lifting the open cover of $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Phi_{\Sigma})$ from Lemma 5.2.
Proof.
Each piece $\widetilde{\mathbf{W}}(P)$ of our Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a Weinstein sector with skeleton $\mathbb{L}(\Sigma/\rho_{P})$ obtained from $T^{*}\widehat{M}_{P}$ for a vertex $P\in\Phi_{\Sigma}$ by the inductive Weinstein handle attachments. Intersections of $\widetilde{\mathbf{W}}(P)$ are Weinstein sectors-with-sectorial-corners whose skeleta are given by intersections of $\mathbb{L}(\Sigma/\rho_{P})$. Since by Lemma 5.2 the union of $\mathbb{L}(\Sigma/\rho_{P})\cong\mathbb{L}(\Sigma/\rho_{P})\times\partial_{\infty}\rho_{P}$ covers $\partial_{\infty}\mathbb{L}(\Sigma)$ when $\rho_{P}$ runs through the rays in $\Sigma$, the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ gives a lift of the open cover of $\partial_{\infty}\mathbb{L}(\Phi_{\Sigma})$ from Lemma 5.2. ∎
Remark 5.8.
The Weinstein domain $\mathbf{W}(\Phi_{\Sigma})$ may be identified with the Weinstein hypersurface $F\subset\partial_{\infty}D$ which gives the sutured Weinstein manifold $(D,F)$ from [GS22, Corollary 6.2.6]. Due to Lemma 3.11(4), $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ gives a Weinstein thickening of $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Sigma)=\widetilde{\mathbb{L}}(\Phi_{\Sigma})$ which is the gluing of local Weinstein thickenings of $\mathbb{L}(\Sigma/\rho_{P})$.